The sum of two numbers is 52 the smaller number is 22 less than the larger number what are the numbers

Answer by ikleyn(52798)

.
smaller + (smaller + 22) = 52


2*smaller = 52 - 22 = 30.


smaller = 30/2 = 15.        ANSWER


greater = 15 + 22 = 37.     ANSWER